your grandmother have you a gold ring as a gift. the gold ring is 5mm thick and has an outside diameter of 12mm and an inside diameter of 11mm. the density of the gold is 1.09 x 10^4 kg/m^3

(a) what is the volume of the ring in cubic millimeter?

(b) what is the mss of the ring in grams?

(c) what is the market value of the ring if the price of gold is P1836.25 per gram?

(d) what is the volume of the ring in m^3?

(e) what is the mass of the ring in kg?


1
Expert's answer
2022-01-14T09:49:02-0500

a)

"V=\pi(d_2^2-d_1^2)h=361~mm^3,"

b)

"m=\rho V=3.94~g,"

c)

"C=Pm=7227.5~peso,"

d)

"V=361\cdot 10^{-9}~m^3,"

e)

"m=3.94\cdot 10^{-3}~kg."